# Which Optimizer Works Best for Physics-Informed Neural Networks and Kolmogorov-Arnold Networks?

### v1

@@ Acknowledgments
4 Appendix A: Lorenz system
@@
Figure 26: Double-Precision PINNs for the Lorenz system: Comparison of PINN predictions using BFGS and SSBroyden with Strong Wolfe line-search against the numerical solution for the Lorenz system. The plots display $x(t)$, $y(t)$, and $z(t)$ over time.
@@
5 Appendix B: Multi-dimensional Rosenbrock function
Considering non-quadratic objective functions Rosenbrock, which is commonly used to evaluate optimization algorithms. The Rosenbrock function is defined as:
@@ 5 Appendix B: Multi-dimensional Rosenbrock function
The SSBroyden method with Strong Wolfe line-search demonstrates performance comparable to BFGS with Strong Wolfe line-search, with the optimization starting from $x_{i}=0.5$. When applied to challenging optimization problems, including the Rosenbrock function—characterized by a narrow, curved valley—and multi-modal landscapes, SSBroyden efficiently converges to the global minimum with accuracy similar to BFGS.
As evident from Figure 27, the number of iterations required to reach the global minimum increases as the dimensionality of the problem grows. However, for the 10-dimensional case, SSBroyden requires more iterations to reach global optimization compared to BFGS. In contrast, for the 2D and 5D cases, both BFGS and SSBroyden require approximately the same number of iterations to achieve the global minimum.
Figure 27: Multi-dimensional Rosenbrock) Convergence of BFGS and SSBroyden with Strong Wolfe line-search for optimizing the Rosenbrock function across different dimensions (2D, 5D, and 10D).
Optimizer, Line-search Optimized parameters Function value Iterations
BFGS with Wolfe 2D $[1.0,1.0]$ $7.07\times 10^{-19}$ 18
SSBroyden with Wolfe 2D $[1.0,1.0]$ $1.56\times 10^{-19}$ 18
BFGS with Wolfe 5D $[1.0,1.0,1.0,1.0,1.0]$ $3.07\times 10^{-20}$ 27
SSBroyden with Wolfe 5D $[1.0,1.0,1.0,1.0,1.0]$ $1.09\times 10^{-21}$ 29
BFGS with Wolfe 10D $[1.0,1.0,...,1.0]$ $4.58\times 10^{-20}$ 44
SSBroyden with Wolfe 10D $[1.0,1.0,...,1.0]$ $6.55\times 10^{-21}$ 61
Table 17: High-dimensional Rosenbrock function Optimization results for the Rosenbrock function across different dimensions using BFGS and SSBroyden with Strong Wolfe line-search. The function value and iteration count show that SSBroyden performs comparably to BFGS in lower dimensions but requires more iterations in the 10D case to achieve a lower function value.

### v2

@@ Acknowledgments
Appendix A Lorenz system
@@
Figure 26: Double-Precision PINNs for the Lorenz system: Comparison of PINNs predictions using BFGS and SSBroyden with Strong Wolfe line-search against the numerical solution for the Lorenz system. The plots display $x(t)$, $y(t)$, and $z(t)$ over time.
@@
Appendix B Multi-dimensional Rosenbrock function
Here, we consider the non-quadratic objective functions Rosenbrock, which is commonly used to evaluate optimization algorithms, defined in two-dimensions as:
@@ Appendix B Multi-dimensional Rosenbrock function
The SSBroyden method with Strong Wolfe line-search demonstrates performance comparable to SSBFGS and BFGS with Strong Wolfe line-search, with the optimization starting from $x_{i}=0.5$. When applied to challenging optimization problems, including the Rosenbrock function—characterized by a narrow, curved valleys and multi-modal landscapes- SSBroyden efficiently converges to the global minimum with accuracy similar to BFGS.
Figure 27 presents a comparison of optimization performance on the Rosenbrock function across four different dimensionalities: 2D, 5D, 10D, and 20D. To ensure a fair comparison, all optimizers were initialized with the same starting point ($\mathbf{x}_{0}=[0.5,\dots,0.5]$), and the stopping criterion was uniformly defined as the gradient norm falling below $10^{-6}$ or a maximum of 5000 iterations. Both methods used the same initial inverse Hessian approximation ($H_{0}=I$) and consistent optimization parameters. The plots highlight the convergence behavior of each method across increasing dimensions, illustrating how the number of iterations and the rate of loss decay vary with problem size. The number of iterations required to reach the global minimum increases as the dimensionality of the problem grows. However, for the 20-dimensional case, SSBFGS and SSBroyden requires more iterations to reach global optimization compared to BFGS. While, for the 2D case, BFGS, SSBFGS and SSBroyden require approximately the same number of iterations to achieve the global minimum.
Figure 27: Convergence comparison of BFGS, SSBroyden, and SSBFGS with Strong Wolfe line-search on the Rosenbrock function across multiple dimensions (2D, 5D, 10D, and 20D).
Dimension Optimizer, Line-search Iterations Optimized parameters Final Loss
2 BFGS with Wolfe 17 [1.0, 1.0] $1.50\times 10^{-15}$
2 SSBroyden with Wolfe 17 [1.0, 1.0] $2.97\times 10^{-15}$
2 SSBFGS with Wolfe 19 [1.0, 1.0] $3.42\times 10^{-20}$
2 Gradient Descent 5000 [0.9689, 0.9386] $9.71\times 10^{-4}$
2 Adam 3899 [1.0, 1.0] $1.24\times 10^{-12}$
5 BFGS with Wolfe 26 [1.0, 1.0, 1.0, 1.0, 1.0] $2.56\times 10^{-17}$
5 SSBroyden with Wolfe 27 [1.0, 1.0, 1.0, 1.0, 1.0] $7.79\times 10^{-17}$
5 SSBFGS with Wolfe 31 [1.0, 1.0, 1.0, 1.0, 1.0] $1.44\times 10^{-17}$
5 Gradient Descent 5000 [0.9970, 0.9939, 0.9879, 0.9759, 0.9522] $7.78\times 10^{-4}$
5 Adam 5000 [1.0, 1.0, 1.0, 1.0, 1.0] $2.77\times 10^{-11}$
10 BFGS with Wolfe 43 [1.0, …, 1.0] $1.53\times 10^{-17}$
10 SSBroyden with Wolfe 57 [1.0, …, 1.0] $1.31\times 10^{-15}$
10 SSBFGS with Wolfe 49 [1.0, …, 1.0] $5.19\times 10^{-17}$
10 Gradient Descent 5000 [0.9999, …, 0.9421] $1.15\times 10^{-3}$
10 Adam 5000 [0.9998, 1.0011, …, 1.0002] $3.77\times 10^{-3}$
20 BFGS with Wolfe 60 [1.0, …, 1.0] $6.35\times 10^{-16}$
20 SSBroyden with Wolfe 81 [1.0, …, 1.0] $1.40\times 10^{-16}$
20 SSBFGS with Wolfe 70 [1.0, …, 1.0] $4.90\times 10^{-16}$
20 Gradient Descent 5000 [1.0, …, 0.9312] $1.64\times 10^{-3}$
20 Adam 5000 [1.0, …, 0.9999] $6.57\times 10^{-7}$
Table 18: Final optimization results for the Rosenbrock function using various methods across dimensions 2D, 5D, 10D, and 20D. All quasi-Newton methods recover the global minimum to machine precision. First-order methods (GD, Adam) either converge slowly or stall at suboptimal values in higher dimensions.